### v2

# A coupled-trajectory approach for decoherence, frustrated hops and internal consistency in surface hopping

Abstract
We address the issues of decoherence, frustrated hops and internal consistency in surface hopping. We demonstrate that moving away from an independent-trajectory picture is the strategy which allows us to propose a robust surface hopping scheme overcoming all these issues at once. Based on the exact factorization and on the idea of coupled trajectories, we consider the swarm of trajectories, that mimic the nuclear dynamics in nonadiabatic processes, as a unique entity. In this way, imposing energy conservation of the swarm and allowing the trajectories to share energy when hops occur clearly indicates the route towards a new surface hopping scheme. Encouraging results are reported, in terms of electronic and vibrational time-dependent properties on the photodynamics of fulvene, modeled with a full-dimensional linear vibronic coupling Hamiltonian.
Nonadiabatic molecular dynamics is an irreplaceable method in the toolkit of physical chemists to model time-resolved photochemistry, the study of the coupled electron-nuclear evolution of molecules upon light excitation[1]. Among the many theories available nowadays, surface hopping, since its first appearance in the literature in 1971[2], has grown to become the method of choice to simulate the nonadiabatic dynamics of molecules and to accompany experimental observations.[3, 4, 5, 6, 7, 8, 9, 10, 11] Despite alternative theories being continuously developed, surface hopping remains very popular thanks to the unrivaled accuracy-cost compromise, to its relative implementation simplicity and especially to the fact that it conveys a very intuitive picture of a nonadiabatic process.[12, 13, 14, 15, 16, 17] In surface hopping, the nuclear wavepacket is represented in terms of a swarm of independent trajectories that classically evolve on the support of electronic potential energy surfaces obtained as the solution of the time-independent Schrödinger equation at a given nuclear configuration. The key nonadiabatic effects, beyond the Born-Oppenheimer approximation, i.e., transitions between electronic states, are mimicked by stochastic hops of the trajectories from one potential energy surface to another usually occurring in localized regions of configuration space. Each trajectory of the swarm is coupled to the quantum-mechanical evolution of the electronic time-dependent state – even though not all surface hopping implementations require this.[18, 19]
Surface hopping has its weaknesses, which have been reported extensively in the literature: it suffers from overcoherence, due to the disconnect between the dynamics of the classical nuclei and of the quantum electrons[20, 21, 22, 23, 24, 25, 16, 26, 27, 28]; when hops occur, energy conservation for the hopping trajectory is enforced, which causes frustrated hops if a trajectory hopping to a state of higher potential energy does not have enough kinetic energy to compensate for the increase in potential[29]; velocity rescaling related to the energy conservation after the hops is applied quite arbitrarily and strongly influence the dynamics;[30, 31, 32, 33, 34] using a swarm of independent trajectories to mimic the evolution of a quantum wavepacket misses non-local quantum-mechanical effects[22, 35, 19, 36, 37]. These weaknesses are consequences of the fact that surface hopping was not deduced originally from first principles, even though derivations [38, 35, 39] and qualitative justifications [40, 41] were proposed later on.
In this work, we address the question: What is missing in surface hopping that would cure all its weaknesses at once? We demonstrate that the idea of coupled trajectories emanating from the exact factorization [42, 43, 44, 45] is capable to indicate the route to follow. Specifically, this route implies that the “surface hopping” idea is used as a tool to simplify conceptually and numerically the evolutions equations derived from the exact factorization, rather than as the working framework. In this way, we do not need to cure the weaknesses of surface hopping, since the starting point is a formulation which is naturally free from these weaknesses. Therefore, we propose a theoretical framework that fully exploits the coupling of the trajectories and that is combined with the algorithmic efficiency of surface hopping, bypassing introducing the issues of decoherence, frustrated hops and internal consistency. Note that a coupled-trajectory-based surface hopping scheme has been developed previously [19, 36], but, as we will show below, it inherits some of the issues of the original surface hopping since the idea of coupled trajectories was not fully “embraced” by all aspects of the algorithm. Instead, our new proposition is robust, stable and accurate, and, thus, we believe that our strategy is an important step forward in the development of novel ideas for nonadiabatic molecular dynamics schemes.
The idea of coupled trajectories emerges naturally in the framework of the exact factorization (EF).[42, 43] When the time-dependent molecular wavefunction is factored as the product of a marginal nuclear amplitude, or wavefunction, and a conditional electronic amplitude, the molecular time-dependent Schrödinger equation (TDSE) can be decomposed as an effective nuclear TDSE and an electronic evolution equation. The nuclear dynamics is generated by time-dependent vector and scalar potentials, reminiscent of classical electromagnetic potentials, representing the effect of the electrons. Formally, an elegant and rigorous procedure can be introduced to derived classical-like Hamilton’s equations from a TDSE, which is the procedure followed in context of EF to deduce classical-like nuclear trajectories.[46] In turn, the nuclei affect the electronic dynamics entering the evolution equation with their momentum/velocity field and with their spatial distribution. As simple as this, the coupled-trajectory idea emerges in the exact factorization: while the momentum field can be interpreted classically, as it is a local property that can be encoded in a trajectory, the very nature of the spatial nuclear distribution is non local. Therefore, to determine the quantum momentum, as we dubbed the term depending on the nuclear distribution, the trajectories become coupled since only the whole ensemble yields the necessary non-local information. Extensive analyses of the time-dependent potentials of the EF have justified additional approximations, namely (i) the use of purely adiabatic forces to drive the nuclear dynamics, since the time-dependent scalar potential reproduces piecewise the adiabatic shapes [40], and (ii) the idea of hopping with associated velocity rescaling, to compensate for the effect of the time-dependent vector potential that couples to the nuclear kinetic energy in the nuclear TDSE [41, 47]. Based on these observations, the coupled-trajectory Tully surface hopping (CT-TSH) algorithm was proposed [19, 36].
CT-TSH is capable to circumvent the issue of overcoherence via the presence of the quantum momentum in the electronic evolution, which encodes, as argued above, information about the nuclear spatial delocalization. Nonetheless, as we will show below and as previously observed [48], CT-TSH suffers from frustrated hops since, apart from the evaluation of the quantum momentum, the trajectories behave as independent entities and their coupling is not fully exploited, in particular to enforce energy conservation after the hops. Instead, by simply understanding that the trajectories need to cooperate to yield a meaningful approximation of the nuclear quantum dynamics, the additional issues discussed above can be easily cured: if the trajectories share their kinetic energy when hops occur, so as to conserve the total energy of the swarm, frustrated hops are avoided (or strongly reduced to a statistically irrelevant amount) and different velocity rescaling schemes do not affect the outcome of the simulation.
Understanding the importance of fully exploiting the coupling among the trajectories in surface hopping leads us to the derivation of the combined coupled-trajectory Tully surface hopping (CCT-TSH), that we illustrate and test in the remainder of this paper. CCT-TSH is based on the idea of “energy sharing” to conserve the energy of the swarm of trajectories and to alleviate the frustrated hops. We propose three schemes for the energy sharing, which are schematically summarized in Fig. 1: (i) the equity-based scheme, (ii) the overlap-based scheme, and (ii) the quantum-momentum-based scheme.
Figure 1: Schematic overview of the CCT-TSH algorithm.
Before detailing the energy-sharing schemes, let us introduce the following quantities. The trajectories are labeled with Greek indices, and the hopping trajectory is the trajectory $\alpha$, with kinetic energy indicated with the symbol $E_{\text{kin}}^{\alpha}$. The variation in potential energy for a hop from state $I$ towards a state $J$ of higher energy is $\Delta E^{\alpha}_{\text{hop}}<0$. The Cartesian components of the vectors are indicated with Latin indices, thus for instance the symbol $v_{i}^{\beta}(t)$ stands for the component $i$ of the velocity of the trajectory $\beta$ at time $t$. For scheme (ii), we define the overlap between two trajectories, i.e., $S_{\alpha\beta}$, as the overlap between two Gaussians centered at the positions of the trajectories $\alpha$ and $\beta$. Note that the algorithm does not depend on the value of the width of the Gaussians, as long as they are the same for all trajectories, since $S_{\alpha\beta}$ is only used to quantify the proximity of the trajectories. For scheme (iii), we identify two channels for energy sharing, one driven by the nonadiabatic couplings between states $I$ and $J$, $\boldsymbol{d}_{IJ}^{\alpha}$, and one driven by the quantum momentum, $\boldsymbol{Q}^{\alpha}$, both evaluated at the position of the hopping trajectory. Therefore, we use the scalar quantities $\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}$ and $\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}$, to decompose the energy variation to be compensated for the hop to occur. Note that these scalar quantities appear in the electronic evolution equation of CT-TSH and CCT-TSH, with $\boldsymbol{f}_{IJ}^{\alpha}$ the time-integrated difference of the adiabatic forces for states $I$ and $J$.
Using the definitions given above, the energy-sharing schemes can be summarized as follow:
  1. (i)
    Equity-based. If the trajectory $\alpha$ has insufficient kinetic energy to compensate a hop from state $I$ to a higher-lying state $J$, the difference $\Delta E_{\text{hop}}^{\text{CCT}}=\Delta E^{\alpha}_{\text{hop}}-E_{\text{kin}}^{\alpha}$ is distributed among all remaining trajectories of the swarm depending on their kinetic energy, but we defined a kinetic energy threshold below which a trajectory is not considered for the energy sharing. If the sum of the kinetic energies of the contributing trajectories is sufficient for the hop, their velocities $v_{i}^{\beta}$ are rescaled as
    ${v_{i}^{\beta}}^{\prime}=v_{i}^{\beta}\sqrt{1+\frac{\Delta E_{\text{hop}}^{\text{CCT}}}{\sum_{\gamma}E_{\text{kin}}^{\gamma}}}$ (1)
    where ${v_{i}^{\beta}}^{\prime}$ is the velocity after rescaling. Weighting the contribution of the trajectories $\beta$ depending on their kinetic energy ensures a minimal impact on the trajectory itself. Since the trajectory $\alpha$ uses its entire kinetic energy for the hop, its new velocity after the hop is $\boldsymbol{v}^{\alpha}=0$.
  2. (ii)
    Overlap-based. If the trajectory $\alpha$ has insufficient kinetic energy to compensate a hop from state $I$ to a higher-lying state $J$, the contribution of the trajectories $\beta$ is determined based on their spatial proximity to the trajectory $\alpha$. The velocity rescaling is done as
    ${v_{i}^{\beta}}^{\prime}=v_{i}^{\beta}\sqrt{1+\frac{\Delta E_{\text{hop}}^{\text{CCT}}S_{\beta\alpha}}{E_{\text{kin}}^{\beta}\sum_{\gamma}S_{\gamma\alpha}}}$ (2)
    where $\sum_{\gamma}S_{\gamma\alpha}$ is only summed over trajectories whose kinetic energy is large enough for the rescaling. The trajectory $\alpha$ uses its entire kinetic energy for the hop, and its new velocity after the hop is $\boldsymbol{v}^{\alpha}=0$.
  3. (iii)
    Quantum-momentum-based. The amount of energy to be rescaled $\Delta E^{\alpha}_{\text{hop}}$ is partitioned in two contributions: One contribution, $\Delta E^{\alpha}_{\text{NAC}}$, that will be compensated by the kinetic energy of the hopping trajectory $\alpha$ and that will be rescaled along the nonadiabatic coupling vector, and a second contribution, $\Delta E^{\alpha}_{\text{Qmom}}$, that will be compensated by the kinetic energy of the whole swarm of trajectories based on (ii). The two contributions are calculated according to
    $\Delta E^{\alpha}_{\text{NAC}}=\Delta E^{\alpha}_{\text{hop}}\frac{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}}{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}+\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}$ (3)
    and
    $\Delta E^{\alpha}_{\text{Qmom}}=\Delta E^{\alpha}_{\text{hop}}\frac{\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}{\boldsymbol{d}_{IJ}^{\alpha}\cdot\boldsymbol{v}^{\alpha}+\boldsymbol{Q}^{\alpha}\cdot\boldsymbol{f}_{IJ}^{\alpha}}$ (4)
    such that $\Delta E^{\alpha}_{\text{hop}}=\Delta E^{\alpha}_{\text{NAC}}+\Delta E^{\alpha}_{\text{Qmom}}$.
In general, in CCT-TSH, we assume that the kinetic energy of the hopping trajectory $\alpha$ is used first, before requesting the cooperation of the surrounding trajectories via the energy sharing. Therefore, the velocity rescaling options used in the equity-based and in the overlap-based sharing schemes are either isotropic, meaning that all components of the velocity vector are equally rescaled, or hybrid, meaning that the component of the velocity along the nonadiabatic coupling vector is rescaled if possible, and otherwise it is isotropic. Nonetheless, we will show that the velocity rescaling procedure does not impact the results.
It is worth discussing here the decoherence problem of surface hopping and its relation to the internal consistency. Internal consistency refers to the agreement between the two ways of estimating the population of the electronic states in surface hopping: the fraction of trajectories associated to, or evolving in, an electronic state (F); the trajectory-averaged population propagated along each trajectory (P), since usually a quantum-dynamical evolution equation for the electronic state is integrated along each trajectory. However, in general and unless decoherence corrections are used, it cannot be ensured that the two estimates agree with each other. The original hopping algorithm [2] ensures this internal consistency by using an average hopping probability over the swarm of trajectories for which the two estimates for the populations are the same. However, as soon as an independent trajectory formalism is invoked and the hopping probability is calculated at the level of a single trajectory, internal consistency is no longer guaranteed. The coupled-trajectory schemes that we propose here are naturally designed to reintroduce the idea of an average hopping probability to restore internal consistency. However, within CT-TSH this might introduce unphysical artifacts. Such an average hopping algorithm will assign the same hopping probability to all trajectories independently of their position. While this might ensure a correct distribution of the wavepacket/swarm with respect to the electronic states, it neglects the locality of the nonadiabatic couplings and the spatial distribution of the wavepacket, causing trajectories to undergo hops even in regions of negligible nonadiabatic coupling. This limitation has previously been overcome through the consensus surface hopping algorithm.[35] For CCT-TSH, we propose, by contrast, to forgo the stochastic hopping algorithm and introduce a deterministic hopping. [49] Since the electronic evolution is derived from the EF equation for the conditional amplitude, we drive the trajectory hops based on the values of the electronic populations predicted by this equation. In a way, the nuclear trajectories in CCT-TSH can be regarded merely as a support on which the time-dependent electronic state and the quantum momentum are evaluated. Therefore, it is crucial that the trajectories evolve following the electronic populations. For this reason, in CCT-TSH, the trajectories evolve in the electronic state with the largest population.
The CCT-TSH algorithms illustrated above are tested on the 30-dimensional linear vibronic coupling (LVC) model of fulvene.[29] Upon photoexcitation to the first excited state S1, fulvene decays through two conical intersections, the first one is a sloped conical intersection seam and is associated to a stretching of the ethenylic bond and the second one is a peaked conical intersection and is associated to the twist of the methylene group [32, 50]. This LVC model of fulvene reproduces only the sloped conical intersection, but not the peaked intersection. However, the latter appears to be not important in the simulated short time dynamics as argued previously [51, 29]. Furthermore, we are interested in correctly capturing the challenging dynamics at the slope conical intersection, where the wavepacket/swarm is partially reflected and thus crosses multiple times the nonadiabatic coupling region. Also, note that the molecular system of fulvene has been shown to be highly sensitive to the parameters of the surface hopping algorithm, in terms of procedure to treat frustrated hops, velocity rescaling and time step.[32, 33, 31] Therefore, it is a suitable test system to verify the robustness of the proposed CCT-TSH schemes. All simulations have been initialized with a swarm of 500 trajectories with positions and momenta Wigner sampled from the vibrational ground state of the electronic LVC ground state potential. The trajectories were propagated with a time step of 0.1 a.t.u.
Figure 2: Adiabatic population decays of S1 predicted by the different surface hopping schemes discussed in the main text. The lighter lines in each panel show the dynamics with different velocity rescaling procedures and treatments of frustrated hops, while the bold darker lines show the average of the individual runs. Solid lines show the average electronic populations (P) and dashed lines the fractions of trajectories (F).
In Fig. 2, we show the adiabatic population curves of the fulvene LVC model upon excitation to the S1 state. For the standard surface hopping approaches, TSH, TSH-ED, CT-TSH, six different dynamics are averaged, using three different schemes to rescale the velocity after a hop, with each option used in combination with velocity inversion or not in the case of frustrated hops. The three schemes for the velocity rescaling are (1) isotropic, (2) along the nonadiabatic coupling vectors, or (3) mixed, meaning that if (2) is not possible then (1). In a standard application of any surface hopping scheme to a molecular problem, each of these approaches would per se be regarded as a justified choice, which is why we chose to use the repetitions with different schemes to give us some information on the uncertainty of the results. It can be clearly seen that for both TSH and CT-TSH internal consistency is not ensured, and in both cases, the electronic populations (P) decay slower than the fractions of trajectories (F). TSH-ED restores, as expected, almost perfect internal consistency. The population dynamics are very similarly predicted by all three methods, with a sharp decrease of S1 population occurring after 250 a.t.u., but since the wavepacket/swarm of trajectories is partially reflected, we observe stepwise decays to around only 10% of S1 population at the end of the simulated dynamics. When looking at CCT-TSH, it can be observed that all three proposed schemes restore closely internal consistency. The population decay predicted using the equity-based and overlap-based energy sharing schemes are almost identical, which points to an encouraging robustness of the algorithm. However, the quantum-momentum-based sharing scheme yields a different population decay with more pronounced coherent oscillations of the population. Given that the quantum-momentum-based energy sharing acts at every surface hop and not only when a hop would be frustrated on the single trajectory level, this result is hardly surprising.
The number of frustrated hops during the dynamics varies strongly between the different methods (see Table 1). As expected, when we allow only the component of the velocity along the nonadiabatic coupling vectors to be rescaled, the number of frustrated hops is quite large in all methods. Applying a decoherence correction as in TSH-ED, not only improves the internal consistency but also strongly decreases the number of frustrated hops. If the full velocity vector can be rescaled isotropically, there are no (or a negligible number of) frustrated hops during the dynamics. The number of frustrated hops during the CT-TSH dynamics is comparable to TSH, but surprisingly much larger if only the component of the velocity along the nonadiabatic coupling vectors can be rescaled. This issue is completely alleviated, albeit by construction, when using the equity-based and overlap-based CCT-TSH schemes. By contrast, the energy sharing scheme based on the quantum momentum shows a large number of frustrated hops. This can be explained by the modified hopping scheme where at each time step the energy sharing is attempted.
Number of frustrated hops
reflect frust hops no reflection
full $v$ mix NACV full $v$ mix NACV
TSH 81 89 490 92 103 591
TSH-ED 0 1 147 0 1 130
CT-TSH 57 119 1121 73 101 2049
Equity Overlap Qmom
CCT-TSH 0 0 138
Table 1: Number of frustrated hops in the different surface hopping dynamics of fulvene.
In order to ascertain that the energy sharing scheme does not lead to artifacts in the spatial distribution of the wavepacket, we report the distribution of the trajectories along the normal mode corresponding to the C=CH2 stretching. This mode was chosen based on previous works where it has been observed that the main sloped conical intersection between S1 and S0 is accessed via a stretching of the C=CH2 bond. [32, 52] Each trajectory was convoluted with a Gaussian of width 0.01 along this mode. We note that in Fig. 3 the $y$-axis is given in mass-weighted normal mode coordinates. Negative values correspond to an elongation of the bond, while positive values correspond to a contraction. The dynamics of TSH, TSH-ED and CT-TSH were averaged over all the six different runs. Over time, the swarm of trajectories shows harmonic oscillations along this bond, initially elongating it, which drives the population decay. After the first elongation, around 500 to 1000 a.t.u., a splitting of the swarm of trajectories can be observed. One part contracts the bond further whereas another part contracts only until the equilibrium position before elongating again. This splitting can be attributed to the electronic population, where the part that decays to the ground state harmonically oscillates between $-4$ and $4$ and the part that gets reflected on the excited state oscillates only in the elongation between 00 and $-4$. In all methods, TSH, TSH-ED and CT-TSH, these two branches are clearly distinguishable between 500 and 2000 a.t.u. A very similar behavior is exhibited by the CCT-TSH trajectories within the equity-based and overlap-based schemes. As with the population traces, their behavior seems to be in excellent agreement with TSH-ED results. The rescaling proportional to the quantum momentum shows a similar overall behavior, however the branching is less clear as the trajectories are overall more spatially delocalized.
Figure 3: Spatial evolution of the swarms of trajectories along the C=CH2 stretching mode. The $y$-axis is given in mass-weighted normal mode coordinates, negative values correspond to a stretching of the C=CH2 bond, positive values to a contraction. TSH, TSH-ED and CT-TSH are averaged over the six different dynamics.
The kinetic energy along this normal mode (shown in Fig. 4) shows that the redistribution of kinetic energy in equity-based and overlap-based CCT-TSH does not alter significantly the average kinetic energy along that normal mode. For TSH, TSH-ED and CT-TSH, the six different runs are shown. The average kinetic energies of all methods are in good agreement and CCT-TSH yields values that lie within the uncertainty spread of the different standard surface hopping schemes. Through the kinetic energy, it becomes clear that the quantum-momentum-based approach does not yield reliable results, as it largely increases the kinetic energy in this mode.
Figure 4: Average kinetic energy of the ensemble of trajectories along the C=CH2 stretching mode.
In summary, we have identified a new and robust strategy to simulate the electron-nuclear dynamics underlying nonadiabatic processes, which is based on the exact factorization and employs the idea of surface hopping to approximate the nuclear dynamics. The crucial ingredient is the coupling of the trajectories, a concept that emerges naturally in the context of the exact factorization, and is the key to achieve internal consistency, to circumvent overcoherence, to alleviate the frustrated hops, and to produce stable results upon changing the velocity rescaling within surface hopping. The coupling of the trajectories is essential to encode non-local, purely quantum mechanical features in the electronic and nuclear dynamics simulated using a surface hopping scheme, to calculate the quantum momentum and to introduce the energy sharing. Three practical schemes were proposed and tested on full-dimensional photodynamics of fulvene, based on a linear vibronic coupling model. Out of three, two schemes were shown to give robust results for electronic and nuclear properties. All the different flavors of CCT-TSH have been implemented in G-CTMQC [53] and are freely available on GitLab, while the LVC fulvene model has been implemented in QuantumModelLib[54].
Acknowledgements.
This work was supported by the French Agence Nationale de la Recherche via the projects Q-DeLight (Grant No. ANR-20-CE29-0014) and STROM (Grant No. ANR-23-ERCC-0002), and by a public grant from the Laboratoire d’Excellence Physics Atoms Light Matter (LabEx PALM) overseen by the French National Research Agency (ANR) as part of the “Investissements d’Avenir” program (reference: ANR-10-LABX-0039-PALM).